# Volumetrically Consistent 3D Gaussian Rasterization

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has enabled photorealistic view synthesis at high inference speeds. However, its splatting-based rendering model makes several approximations to the rendering equation, reducing physical accuracy. We show that splatting and its approximations are unnecessary, even within a rasterizer; we instead volumetrically integrate 3D Gaussians directly to compute the transmittance across them analytically. We use this analytic transmittance to derive more physically-accurate alpha values than 3DGS, which can directly be used within their framework. The result is a method that more closely follows the volume rendering equation (similar to ray-tracing) while enjoying the speed benefits of rasterization. Our method represents opaque surfaces with higher accuracy and fewer points than 3DGS. This enables it to outperform 3DGS for view synthesis (measured in SSIM and LPIPS). Being volumetrically consistent also enables our method to work out of the box for tomography. We match the state-of-the-art 3DGS-based tomography method with fewer points.
@@ 1 Introduction
Recently, there has been tremendous progress in view synthesis methods using differentiable volume rendering [35, 25, 37, 49, 5]. These techniques can be broadly categorized as rasterization or ray-tracing methods. Both produce photorealistic results but with distinct trade-offs. Rasterization-based methods are faster for both training and inference, but cannot render the complex camera effects that ray-tracing methods can, like lens blur and distortion.
@@
Recent works have taken inspiration from ray-tracing-based methods to make 3DGS more physically-accurate [22, 51, 24]. However, like 3DGS, they all rely on splatted 2D Gaussians for color computation instead of actually volumterically integrating the 3D Gaussian.
We identify that we do not need splatting, including all its approximations, to have a fast rasterization-based technique. Instead of splatting, we analytically evaluate the volume rendering equation by integrating Gaussians directly in 3D space (assuming correct sorting and no overlap) — all within a rasterizer. See Fig. 1 for an illustration.
@@
In practice, our method consistently matches or exceeds 3DGS’s view synthesis quality (measured by LPIPS and SSIM) over a wide variety of scenes (Sec. 6.1). Our method also works out-of-the-box for 3DGS-based tomographic reconstruction and matches a recent state-of-the-art (SOTA) method’s quality with a lower memory footprint (Sec. 6.2).
@@ 2.1 Volume rendering with ray marching
Image-based rendering (view synthesis) works have a long history in computer graphics [17, 30, 43, 9]. Recently, NeRF and its followups [35, 49, 3, 15, 37] pioneered using differentiable volume rendering [34] and have achieved photorealistic reconstruction. These methods use neural networks and grids to represent a scene’s volumetric density and color. Colors along rays are computed by evaluating the volume rendering equation using ray marching. Ray marching typically requires several samples along the ray, each of which involves expensive neural network queries making the (differentiable) rendering process slow. There has been work on real-time NeRF rendering approaches [13, 41, 10, 42, 47, 46] but these are typically multi-stage approaches, and thus overall still slow.
@@ 2.2 Rasterizing point-based volumes
Rasterization-based methods [29, 25, 58] instead represent the scene using explicit primitives, like 3D Gaussians, which are projected onto screen space and composited together. Tile-based rasterizers like 3DGS are fast because they exploit the coherence across pixels within the tile by making some assumptions: a) primitives don’t overlap; b) sorting order for all primitives is consistent across all pixels in a tile. 3DGS also assumes that blending weights (volumetric transmittance) can be approximated using 2D Gaussians, a result of splatting 3D Gaussians onto the screen [57].
@@
While these methods eliminate the affine approximation, none of them compute physically-accurate transmittance. Our method, by analytically integrating 3D Gaussians (instead of splatting them onto the screen) can do so, resulting in a more physically-based rendering model. Hamdi et al. [19] replace 3D Gaussians with generalized exponentials which are then splatted onto the screen; their approach makes primitives more opaque, resulting in a more compressed representation.
Radl et al. [39] reduce popping artifacts due to inconsistent primitive sorting and Hou et al. [21] sidestep primitive sorting altogether by using order-independent transparency. Many recent works [26, 48, 32, 27, 6] have improved the densification heuristics in 3DGS and reduced its dependency on initialization. Our method is a drop-in replacement for the alpha computation in 3DGS, making it compatible with all these recent works as well.
@@ 2.3 Ray-tracing point-based volumes
Recently, there have been several concurrent works that ray-trace point-based volumes for view synthesis [36, 11, 31]. These methods vary in their physical-correctness: a) Moenne-Loccoz et al. [36] approximate the volume rendering equation by placing samples at locations of maximum contribution for each primitive along the ray; b) Condor et al. [11], like us, derive a closed-form expression for transmittance across 3D Gaussian and Epanechnikov kernels; c) Mai et al. [31] compute the volume rendering equation exactly (including sorting and overlap handling) for constant density ellipsoids. These methods, however, are much slower to train and render than rasterization-based ones and require specialized hardware for fast ray-tracing.
@@ 2.4 Computed Tomography
Tomographic reconstruction aims to estimate the underlying 3D density of an object from 2D projections. Classical analytic methods like FDK [14] invert the radon transform [40] and are fast but struggle with sparse measurements. Recently, NeRF-based approaches [54, 52, 53, 8, 44] have shown promising results, but are slow due to ray marching. Several 3DGS-based tomography approaches have been proposed [7, 16, 55]. Zha et al. [55] are the state-of-the art and achieve high quality by fixing integration bias in 3DGS’s Splatting routine.
@@ 3.2 Volume rendering of 3D Gaussians
We briefly recap the Elliptical Weighted Average (EWA) splatting algorithm [58] used by 3DGS [25] and point out the different approximations it makes by underlining them.
The density field
@@
which are assumed to not overlap, have view-dependent color , and be sorted front-to-back [58]. Our method adopts these assumptions too but relaxes the ones that follow, which are necessary for splatting. To splat the 3D Gaussians onto the image plane, EWA (and 3DGS), linearizes the exponential transmittance () and assume Gaussians are not self-occluding, resulting in the alpha blending equation describing the color at pixel
@@
is the Jacobian of the affine approximation to the camera projection transform. For more details please see Zwicker et al. [58]. We stress that the values are approximate, even when the sorting and non-overlapping assumptions are met, due to the three additional approximations made by splatting. Our method does not perform splatting, thereby relaxing the extra approximations in 3DGS. Instead we compute by analytically evaluating Eq. 2, which accumulates the density along ray . This results in an efficient rendering method that matches or exceeds SOTA results in both view synthesis and sparse-view tomography.
@@ 4.1 Alpha blending without splatting
In this section, we describe how our method, which analytically integrates transmittance, can also be written as alpha blending due to the properties of exponential transport. Substituting Eq. 3 into Eq. 1 results in the color for a pixel (with the ray passing through its center) being
Here are the limits of Gaussian along the ray.
Assuming the Gaussians are sorted (front to back) and non-overlapping, we can separate out the transmittance due to the previous Gaussians
@@
We observe that this equation is identical to the alpha-blending equation, Eq. 5, adopted by 3DGS. The difference is that Eq. 14 requires accurately evaluating , the accumulated density of the -th primitive. Note that thus far, we do not rely on any specific property of 3D Gaussians; we only require the primitives to be non-overlapping. This opens up a design space for other compact kernels which we leave for future work; we use 3D Gaussians. In the next section, we show that the values in Eq. 14 can be computed analytically for 3D Gaussians. The result closely matches ray tracing results (assuming no overlap and correct sorting), but uses rasterization for efficient rendering, see Appendix A.
@@ 4.2 Analytic transmittance computation
where is a shorthand for , the maximum value of the Gaussian along the ray. The 1D Gaussian has mean and variance (proof in Appendix B)
@@ 4.3 Benefits of accurate alpha computation
Comparing our alpha derivation to 3DGS’s, we note that 3DGS makes three extra approximations (see Sec. 3.2): it assumes a) linearization of the exponential; b) no-self occlusion; c) linearization of the covariance. Each of these makes renderings less physically based and hurt view synthesis quality.
The result of the first two approximations is that for 3DGS is given by a 2D Gaussian, see Eq. 6. This 2D Gaussian only reaches its maximum opacity of 1 at its center (Fig. 2 (b), dotted), limiting 3DGS’s ability to represent opaque surfaces (). As a consequence, our method is better at fitting piecewise constant textures than 3DGS for a fixed budget of Gaussians. We demonstrate this through examples in and Fig. 2 and Fig. 3. In Fig. 2 we plot the opacity, , of a volumetrically rendered Gaussian with varying density and scale. We observe that when increasing the density , our method results in opacity close to 1 over a much larger region (Fig. 2 (b), solid), thereby enabling it to better represent opaque surfaces (Fig. 2, bottom row).
Another consequence of splatting a 3D primitive to 2D is that it discards the z-scale (along the camera ray) of the Gaussian, which makes renderings invariant to it (Fig. 2 (a), dotted). This is physically inaccurate; larger z-scales should increase a Gaussian’s contribution, making it more opaque. Our method captures this effect, as shown in Fig. 2 (a) (solid). Finally, the linearization of the covariance results in artifacts away from the point of linearization (the center of the Gaussian), as has been shown by several prior works [24, 56, 12, 18, 50, 23]. Since our method directly integrates Gaussians in ray-space, it doesn’t have this issue and the artifacts it causes.
The overall effect is that our method can better fit opaque objects. We demonstrate this in Fig. 3 where we have fit 3D Gaussians to a single image (Fig. 3(a)). We observe that our method (Fig. 3(b)) better approximates sharp edges and constant regions versus 3DGS (Fig. 3(c)), which blurs edges and shows bleeding artifacts.
@@ 5 Implementation details
We make a few small changes to 3DGS’s adaptive density control process for our method: a) we halve density after splitting and cloning; b) we additionally also split points with high density ; c) we prune points with low density (based on i.e. density before reparameterization); d) we disable opacity reset since we found it didn’t help. We list other hyperparameters and implementation details in Appendix C.
@@ 6.1 View synthesis
We evaluate our method on MipNeRF-360, DeepBlending and Tanks and Temples&Datasets. For DeepBlending and Tanks&Temples we use the same scenes as 3DGS. We compare with 3 kinds of baselines, see Tab. 1. NeRF and voxel-based methods [15, 37, 4], point-based ray-tracing methods [31, 11], and rasterization-based 3DGS follow-ups that change the alpha computation similar to what we do [19]. There are several works for improving 3DGS along various axes [26, 27] which are orthogonal to our contribution. We only show comparisons with methods that directly change the alpha computation and are rasterization-based. We will add comparisons for ray tracing-based methods once their code is released.
@@
We consistently match or outperform rasterization-based methods such as 3DGS and GES on SSIM and LPIPS. Since SSIM and LPIPS are more sensitive to edge quality and perceptual similarity, our method performs well on these metrics compared to PSNR, where the ability of 3DGS to render more diffuse textures skews results in its favor. Our method allows 3D Gaussian primitives to represent opaque textures better than 3DGS, leading to comparable or better quality. While GES produces fewer primitives, ours has better reconstruction quality and is compatible with the primitives proposed in their work. For point-based ray-tracing methods, we have better reconstruction performance than Condor et al. [11], which is closest to our work algorithmically. EVER [31] has better reconstruction quality for PSNR and SSIM. But since it uses ray tracing, it has a much lower FPS and higher training times than ours. We report per scene reconstruction metrics for our method in Appendix E.
@@
Since our method is rasterization-based, we are faster than all the ray tracing-based methods for rendering images (and similar to 3DGS / GES). While GES has the lowest storage requirements and highest FPS (though its reconstruction quality is worse than ours). Our training times are much lower than point-based ray-tracing methods and NeRF-based methods, but slightly slower than 3DGS. We use a Nvidia 3090 Ti for training and evaluation of both our method and 3DGS. For MipNeRF-360 scenes, on average 3DGS and our method generate 2.62M and 2.93M points respectively, and the training times for 30000 iterations are 51.5 min and 61 min respectively. Similarly for Tanks&Temples, on average 3DGS and our method generate 1.79M and 1.72M points respectively, and the training times for 30000 iterations are 20.5 min and 33.2 min respectively. The slowdown in training is due to slightly more involved computations in the forward and backward pass, and extra computations in the vertex shading phase for our method compared to 3DGS; more details are in Appendix D.
Reconstruction quality across memory budgets:
In Fig. 5, we compare the performance of our method against 3DGS across different memory budgets, i.e. across a different number of maximum primitives that each method is allowed to generate. We stop densification and pruning once the optimization reaches the maximum number of allowed primitives. For each memory budget, we average the results over 4 scenes, kitchen, stump, train, counter. For the same memory budget, our method consistently outperforms 3DGS on both SSIM and LPIPS. This holds across a range of memory budgets as shown in Fig. 5. Recall that our method represents opaque textures better than 3DGS for the same number of primitives, as demonstrated in Fig. 3 and Fig. 2. Since LPIPS and SSIM measure perceptual similarity and edge quality, this also translates to quantitative improvements as demonstrated in Fig. 5.
| PSNR | SSIM | LPIPS | Points | |
| 3DGS | 21.38 | 0.588 | 0.360 | 3.4M |
| 3DGS + Our HyperParams | 20.66 | 0.571 | 0.329 | 7.4M |
| Ours w/o Reparam | 21.23 | 0.600 | 0.309 | 7M |
| Ours w EVER Reparam | 20.04 | 0.519 | 0.396 | 9.4M |
| Ours | 21.41 | 0.615 | 0.306 | 3.8M |
Ablation study:
We conduct an ablation study to assess the impact of hyperparameters and density parameterization on our method, with results shown in Tab. 2. We compare our method with 3DGS (row 1), and for fairness, we also run 3DGS with the hyperparameters used by our method. Using 3DGS with our hyper-parameters (row 2) improves LPIPS, but worsens PSNR and SSIM while generating almost 2x more primitives compared to the base 3DGS configuration (row 1). Our method (row 4) fares relatively better on all metrics. We also run our method without density reparameterization (row 3); we ensure positive density by applying the softplus activation. We observe improvement in LPIPS and SSIM compared to 3DGS but at almost double the number of primitives. We also experiment with the parameterization proposed in Mai et al. [31] (row 4), which results in poorer SSIM and LPIPS while producing too many points compared to our method (row 5).

### v3

@@ Abstract
Recently, 3D Gaussian Splatting (3DGS) has enabled photorealistic view synthesis at high inference speeds.
However, its splatting-based rendering model makes several approximations to the rendering equation, reducing physical accuracy.
We show that the core approximations in splatting are unnecessary, even within a rasterizer;
we instead volumetrically integrate 3D Gaussians directly to compute the transmittance across them analytically.
We use this analytic transmittance to derive more physically-accurate alpha values than 3DGS, which can directly be used within their framework.
The result is a method that more closely follows the volume rendering equation (similar to ray-tracing) while enjoying the speed benefits of rasterization.
Our method represents opaque surfaces with higher accuracy and fewer points than 3DGS.
This enables it to outperform 3DGS for view synthesis (measured in SSIM and LPIPS).
Being volumetrically consistent also enables our method to work out of the box for tomography. We match the state-of-the-art 3DGS-based tomography method with fewer points. Our code is publicly available at:
https://github.com/chinmay0301ucsd/Vol3DGS
@@ 1 Introduction
Recently, there has been tremendous progress in view synthesis methods using differentiable volume rendering [38, 28, 40, 52, 5]. These techniques can be broadly categorized as rasterization or ray-tracing methods. Both produce photorealistic results but with distinct trade-offs.
@@
Recent works have taken inspiration from ray-tracing-based methods to make 3DGS more physically-accurate [24, 54, 26]. However, these methods don’t volumetrically integrate 3D Gaussians for color computation. Instead, like 3DGS, they rely on splatting, i.e., projecting 3D Gaussians to 2D Gaussians in screen space.
We identify that we do not need splatting, including all its approximations, to have a fast rasterization-based technique. Instead of splatting, we analytically evaluate the volume rendering equation by integrating Gaussians directly in 3D space (assuming correct sorting and no overlap) — all within a rasterizer. See Fig. 1 for a qualitative illustration.
@@
In practice, our method consistently matches or exceeds 3DGS’s view synthesis quality (as measured by LPIPS and SSIM) over a wide variety of scenes (Sec. 6.1).
Our method also works out-of-the-box for 3DGS-based tomographic reconstruction and is able to match a recent state-of-the-art method’s quality with a lower memory footprint (Sec. 6.2).
@@ 2.1 Volume rendering with ray marching
Image-based rendering (view synthesis) works have a long history in computer graphics [19, 33, 46, 11]. Recently, NeRF and its followups [38, 52, 3, 17, 40] pioneered using differentiable volume rendering [37] and have achieved photorealistic reconstruction. These methods use neural networks and grids to represent a scene’s volumetric density and color. Colors along rays are computed evaluating the volume rendering equation using ray marching. Ray marching typically requires several samples along the ray, each of which involves expensive neural network queries making the (differentiable) rendering process slow. There has been work on real-time NeRF rendering approaches [15, 44, 12, 45, 50, 49] but these are typically multi-stage approaches, and thus overall still slow.
@@ 2.2 Rasterizing point-based volumes
Rasterization-based methods [32, 28, 61] instead represent the scene using explicit primitives, like 3D Gaussians, which are projected onto screen space and composited together. Tile-based rasterizers like 3DGS are fast, because they exploit the coherence across pixels within the tile by making some assumptions: a) primitives don’t overlap; b) sorting order for all primitives is consistent across all pixels in a tile. 3DGS also assumes that blending weights (volumetric transmittance) can be approximated using 2D Gaussians, a result of splatting 3D Gaussians onto the screen [60].
@@
While these methods eliminate the affine approximation, none of them compute physically-accurate transmittance. Our method, by analytically integrating 3D Gaussians (instead of splatting them onto the screen) is able to do so, resulting in a more physically-based rendering model. Hamdi et al. [21] replace 3D Gaussians with generalized exponentials which are then splat onto the screen; their approach enables primitives to be more opaque, resulting in a more compressed representation. Radl et al. [42] reduce popping artifacts due to inconsistent primitive sorting and Hou et al. [23] sidestep primitive sorting altogether by using order-independent transparency. Many recent works [29, 51, 35, 30, 8] have improved the densification heuristics in 3DGS and reduced its dependency on initialization. Our method is a drop-in replacement for the alpha computation in 3DGS, making it compatible with all these recent works as well.
@@ 2.3 Ray-tracing point-based volumes
Recently, there have been several concurrent works that ray-trace point-based volumes for view synthesis [39, 13, 34]. These methods vary in their physical-correctness: a) Moenne-Loccoz et al. [39] approximate the volume rendering equation by placing samples at locations of maximum contribution for each primitive along the ray; b) Condor et al. [13], like us, derive a closed-form expression for transmittance across 3D Gaussian and Epanechnikov kernels; c) Mai et al. [34] compute the volume rendering equation exactly (including sorting and overlap handling) for constant density ellipsoids. These methods, however, are much slower to train and render than rasterization-based ones, and require specialized hardware for fast ray-tracing.
@@ 2.4 Computed Tomography
Tomographic reconstruction aims to estimate the underlying 3D density of an object from 2D projections. Classical analytic methods like FDK [16] invert the Radon transform [43] and are fast but struggle with sparse measurements.
Recently, NeRF-based approaches [57, 55, 56, 10, 47] have shown promising results, but are slow due to ray marching. Several 3DGS-based tomography approaches have been proposed [9, 18, 58]. Zha et al. [58] are the state-of-the art and achieve high quality by fixing integration bias in 3DGS’s Splatting routine.
@@ 3.2 Volume rendering of 3D Gaussians
We briefly recap the Elliptical Weighted Average (EWA) splatting algorithm [61] used by 3DGS [28], and point out the different approximations it makes by underlining them. The density field
@@
which are assumed to not overlap, have view-dependent color , and be sorted front-to-back [61]. We also adopt these two assumptions. To compute pixel color, both EWA and 3DGS, linearize the exponential transmittance (), an assumption not strictly necessary for splatting. Our method relaxes this assumption and subsequent assumptions required for splatting. For splatting, both EWA and 3DGS assume Gaussians are not self-occluding, resulting in the alpha blending equation describing the color at pixel
@@
is the affine approximation (Jacobian) to the camera projection matrix. For more details please see Zwicker et al. [61]. We stress that the values are approximate, even when the sorting and non-overlapping assumptions are met, due to the three additional approximations made by splatting. We do not perform splatting, thereby relaxing the extra approximations 3DGS entails. Instead we compute by analytically evaluating Eq. 2, which accumulates the density along ray . This results in an efficient rendering method that matches or exceeds state-of-the-art results in both view synthesis and sparse-view tomography.
@@ 4.1 Alpha blending without splatting
In this section we describe how our method, which analytically integrates transmittance, can also be written as alpha blending due to the properties of exponential transport. Substituting Eq. 3 into Eq. 1 results in the color for a pixel (with the ray passing through its center) being
Here are the limits of Gaussian along the ray. Assuming the Gaussians are sorted (front to back) and non-overlapping, we can separate out the transmittance due to the previous Gaussians
@@
We observe that this equation is identical to the alpha-blending equation, Eq. 5, adopted by 3DGS. The difference is that Eq. 14 requires accurately evaluating , the accumulated density of the -th primitive. Note that thus far, we do not rely on any specific property of 3D Gaussians; we only require the primitives to be non-overlapping. This opens up a design space for other compact kernels which we leave for future work; we use 3D Gaussians. In the next section, we show that the values in Eq. 14 can be computed analytically for 3D Gaussians. The result closely matches ray tracing results (assuming no overlap and correct sorting), but uses rasterization for efficient rendering, see Supp.A.
@@ 4.2 Analytic transmittance computation
where is the maximum value of the Gaussian along the ray, and the 1D Gaussian has mean and variance (proof in Supp.B)
@@ 4.3 Benefits of accurate alpha computation
Comparing our alpha derivation to 3DGS’s, we note that 3DGS makes three extra approximations (see Sec. 3.2): it assumes a) linearization of the exponential; b) no-self occlusion; c) linearization of the covariance. Each of these make renderings less physically-based and hurt view synthesis quality.
The result of the first two approximations is that for 3DGS is given by a 2D Gaussian, see Eq. 6. This 2D Gaussian only reaches its maximum opacity of 1 at its center (Fig. 2 (b), dotted), limiting 3DGS’s ability to represent opaque surfaces (). As a consequence, our method is better at fitting piecewise constant textures than 3DGS for a fixed budget of Gaussians. We demonstrate this through examples in Fig. 2 and Fig. 3. In Fig. 2 we plot the opacity, , of a volumetrically rendered Gaussian with varying density and scale. We observe that when increasing the density , our method results in opacity close to 1 over a much larger region (Fig. 2 (b), solid), thereby enabling it to better represent opaque surfaces (Fig. 2, bottom row). Another consequence of splatting a 3D primitive to 2D is that it discards the z-scale (along the camera ray) of the Gaussian, which makes renderings invariant to it (Fig. 2 (a), dotted). This is physically inaccurate; larger z-scales should increase a Gaussian’s contribution, making it more opaque. Our method captures this effect, as shown in Fig. 2 (a) (solid). Finally, the linearization of the covariance results in artifacts away from the point of linearization (the center of the Gaussian), as has been shown by several prior works [27, 59, 14, 20, 53, 25]. Since our method directly integrates Gaussians in ray-space, it doesn’t have this issue and the artifacts it causes.
The overall effect is that our method can better fit opaque objects. We demonstrate this in Fig. 3 where we have fit 3D Gaussians to a single image (Fig. 3(a)). We observe that our method (Fig. 3(b)) better approximates sharp edges and constant regions versus 3DGS (Fig. 3(c)), which blurs edges and shows bleeding artifacts. We also validate this observation in a synthetic multiview setting in Supp.C.
@@ 5 Implementation details
We make a few small changes to 3DGS’s adaptive densification process for our method: a) we halve density after splitting and cloning; b) we also split points with high density ; c) we prune points with low density (based on i.e. density before reparameterization); d) we disable opacity reset since we found it didn’t help; We list other hyperparameters and implementation details in Supp.D.
@@ 6.1 View synthesis
We evaluate our method on MipNeRF-360, DeepBlending, and Tanks&Temples Datasets. For DeepBlending and Tanks&Temples we use the same scenes as 3DGS. We compare with two kinds of baselines. NeRF and voxel-based methods [17, 40, 4, 6], and point-based methods (rows below Zip-NeRF in Tab. 1) that include ray-tracing [34, 13], and rasterization-based 3DGS follow-ups that change the alpha computation similar to what we do [21]. Many works improve 3DGS along various axes [29, 30] which are orthogonal to our contribution. We only show comparisons with methods that directly change the alpha computation and are rasterization-based.
@@
We consistently match or outperform rasterization-based methods such as 3DGS and GES on SSIM and LPIPS. Since SSIM and LPIPS are more sensitive to edge quality and perceptual similarity, our method performs well on these metrics compared to PSNR, where the ability of 3DGS to render more diffuse textures skews results in its favor. Our method allows 3D Gaussian primitives to represent opaque textures better than 3DGS, leading to comparable or better quality. While GES produces fewer primitives, ours has better reconstruction quality and is compatible with the primitives proposed in their work. For point-based ray-tracing methods, we have better reconstruction performance than Condor et al. [13], which is closest to our work algorithmically. EVER [34] has better reconstruction quality, but since it uses ray tracing, it has a much lower FPS and higher training times than ours. ZipNeRF [6] achieves state-of-the-art performance but is prohibitively slow to train and render. We report per-scene metrics from our method on all three datasets in Supp.E.
@@
Since our method is rasterization-based, we are faster than all the ray tracing-based methods for rendering images (and similar to 3DGS / GES). While GES has the lowest storage requirements and highest FPS (though its reconstruction quality is worse than ours). Our training times are much lower than point-based ray-tracing and NeRF-based methods, but slightly slower than 3DGS, see Supp.F. The training slowdown is due to more complex computations in the forward and backward pass, and extra computations in the vertex shading phase for our method compared to 3DGS; more details are in Supp.F. We also include details on our ZipNeRF training setup in Supp.F. We show that our method outperforms 3DGS across a wide range of memory budgets in Supp.G. We ablate our method and present the results in Supp.H.